# HVAC-EAR: EAVESDROPPING HUMAN SPEECH USING HVAC SYSTEMS

### v1

@@ Abstract
Pressure sensors are widely integrated into modern Heating, Ventilation and Air Conditioning (HVAC) systems. As they are sensitive to acoustic pressure, they can be a source of eavesdropping. This paper introduces HVAC-EAR, which reconstructs intelligible speech from low-resolution, noisy pressure data with two key contributions: (i) We achieve intelligible reconstruction from as low as 0.5 kHz sampling rate, surpassing prior work limited to hot word detection, by employing a complex-valued conformer with a Complex Unified Attention Block to capture phoneme dependencies; (ii) HVAC-EAR mitigates transient HVAC noise by reconstructing both magnitude and phase of missing frequencies. For the first time, evaluations on real-world HVAC deployments show significant intelligibility, raising novel privacy concerns.
@@ 1 Introduction
ii) Transient noise: To resist transient HVAC noise, HVAC-EAR jointly reconstructs clean magnitude and phase from aliased components using a complex-valued network. Unlike prior real-valued approaches [5, 6, 7], HVAC-EAR leverages complex spectrograms and a complex multi-resolution STFT loss to recover intelligible speech with clean phases critical for enhancement [15] (see Section 4.4).

### v2

@@ Abstract
Pressure sensors are widely integrated into modern Heating, Ventilation and Air Conditioning (HVAC) systems. As they are sensitive to acoustic pressure, they can be a source of eavesdropping. We introduce HVAC-EAR, which reconstructs intelligible speech from low-resolution, noisy pressure data with two key contributions: (i) We achieve intelligible reconstruction from as low as 0.5 kHz sampling rate, surpassing prior work limited to hot word detection, by employing a complex-valued conformer with a Complex Unified Attention Block to capture phoneme dependencies; (ii) We mitigate transient HVAC noise by reconstructing both magnitude and phase of missing frequencies. For the first time, evaluations on real-world HVAC deployments show significant intelligibility up to 1.2 m distance, raising novel privacy concerns.
@@ 1 Introduction
ii) Transient noise: To resist transient HVAC noise, HVAC-EAR jointly reconstructs clean magnitude and phase from aliased components using a complex-valued network. Unlike prior real-valued approaches [5, 6, 7], HVAC-EAR leverages complex spectrograms and a complex multi-resolution STFT loss to recover intelligible speech with clean phases critical for enhancement [15] up to 1.2 m (see Section 4.4).